# A Competitive Showcase of Quantum versus Classical Algorithms in Energy Coalition Formation

### v1

@@ Abstract
The formation of energy communities is pivotal for advancing decentralized and sustainable energy management. Within this context, Coalition Structure Generation (CSG) emerges as a promising framework. The complexity of CSG grows rapidly with the number of agents, making classical solvers impractical for even moderate sizes (number of agents>30). Therefore, the development of advanced computational methods is essential. Motivated by this challenge, this study conducts a benchmark comparing classical solvers with quantum annealing on Dwave hardware and the Quantum Approximation Optimization Algorithm (QAOA) on both simulator and IBMQ hardware to address energy community formation. Our classical solvers include Tabu search, simulated annealing, and an exact classical solver. Our findings reveal that Dwave surpasses QAOA on hardware in terms of solution quality. Remarkably, QAOA demonstrates comparable runtime scaling with Dwave, albeit with a significantly larger prefactor. Notably, Dwave exhibits competitive performance compared to the classical solvers, achieving solutions of equal quality with more favorable runtime scaling.
@@ I Introduction
To explore the potential of quantum computing in optimization problems it is encouraging to benchmark quantum solvers against classical solvers in tackling problems that are sufficiently hard and exceed the capabilities of state-of-the-art classical solvers. This exploration becomes even more intriguing when problems get already hard in intermediate sizes where existing quantum hardware can be effectively leveraged. It is also crucial that the problem can be formulated compatible with existing quantum algorithms and hardware Abbas et al. (2023). As we explain next, energy coalition formation applying the Coalition Structure Generation (CSG) game stands out as an exemplary candidate for such benchmarking, given its computational complexity and profound real-world significance within the industry. The formation of energy communities represents a significant step towards decentralized and sustainable energy management, offering opportunities for local empowerment, renewable energy adoption, and enhanced resilience in energy systems Tushar et al. (2020); Luo and Liu (2021); Blenninger et al. (2024, 2024).
@@
CSG provides a framework for understanding how agents, can form coalitions denoted by $C_{l}$ (with $l$ being the index of the coalition) to optimize their collective benefits Bandeiras et al. (2023); Moafi et al. (2023); Han et al. (2018); Rahwan et al. (2015). A coalition within this context can be characterized by a characteristic function, typically denoted as $v(C_{l})$, which assigns values to different coalitions. This specific type of cooperative game, defined by the characteristic function, is commonly referred to as a characteristic function game (CFG). The value function provides a measure of the cooperative advantage or payoff that the agents in the coalition can attain by working together, as opposed to operating independently Bachrach et al. (2013); Deng and Papadimitriou (1994).
@@
The task of creating coalition structures is highly challenging, especially given the presence of $n$ agents, leading to a staggering $O(n^{n})$ possible partitions. Currently, the only algorithm capable of finding an optimal solution is the Dynamic Programming algorithm, achieving this in $O(3^{n})$ time Rahwan and Jennings (2008). This approach is limited to systems consisting of a few tens of agents ( i.e. $n<$ 25). CFG can be transformed into an approximately equivalent game represented using the induced subgraph game (ISG) Bistaffa et al. (2022). In this representation, nodes correspond to agents, and coalition values are reflected in pairwise interactions within a weighted graph, such that $v\left(C_{l}\right)=\sum_{(i,j)\in C_{l}}w_{ij}$ Deng and Papadimitriou (1994); Bachrach et al. (2013). The transformation of CFG to ISG allows the problem to be recast as quadratic unconstrained binary optimization (QUBO), enabling more computationally efficient approaches. Nonetheless, it is crucial to note that the problem remains NP-complete Bachrach et al. (2013) such that leading classical solvers, such as DyCE Bistaffa et al. (2014) and CFSS Bistaffa et al. (2014) are limited in scalability and memory requirements. For instance, DyCE Bistaffa et al. (2014), which employs dynamic programming, is restricted to systems with around 25 agents due to its exponential memory complexity ($O\left(2^{n}\right)$) and lacks an anytime approach. CFSS can handle larger but shallow problems, yet still explores all possible solutions for complete graphs with a complexity of $O(n^{n})$.
@@
Figure 1: (a) Schematic representation of the workflow: The Distribution System Operator (DSO) strategically evaluates the potential of prosumer communities, aiming to minimize network management costs through optimized power flow analysis. The DSO’s decision-making involves partitioning prosumers into net-metered coalitions, treating it as a Characteristic Function Game (CFG) problem to optimize overall network efficiency. We transform the CFG formulation into the Induced Subgraph Game (ISG). An approximate solution to the ISG is found through iterative splitting coalitions to bipartitions until no value-increasing bipartitions remain. In the benchmarking phase, the efficiency of both quantum and classical solvers is assessed in the iterative splitting process.
In this work, we propose a novel formulation of energy coalition formation that can be encapsulated within the framework of CFG. Our formulation specifically addresses the value that can be created by cooperative distributed energy resources utilization by net-metered energy coalitions, as well as the costs in terms of network constraint management that can occur if these coalitions are too dispersed. Additionally, we introduce a novel method for transforming the CFG framework into an ISG. We then apply the approximate method introduced in Venkatesh et al. (2023a) to optimize energy community formation.
Under this approximation, the complexity of an exact solver remains $O(n^{3})$ in time while numerical observation shows that an approximate classical solver can solve the problem in polynomial time. Our performance evaluation involves benchmarking an exact classical solver, Tabu search, simulated annealing, and QB-solve (an approximate classical solver developed by Dwave that employs the Tabu algorithm Booth et al. (2017)) against quantum annealing on Dwave hardware, and the Quantum Approximation Optimization Algorithm (QAOA) on both simulators and IBMQ hardware.
Encouragingly, Dwave shows a competitive performance compared to our approximate classical solver. While it achieves solutions of equal quality to those found by the classical solvers, its runtime scales more favorably. This contrasts with many previous studies that demonstrate the poor performance of Dwave compared to state-of-the-art classical solvers on fully connected problems Mohseni et al. (2022); Au-Yeung et al. (2023); Vert et al. (2021); Carugno et al. (2022). In previous works, the better performance of Dwave in comparison with classical solvers is mostly shown for the class of logical-planted problems that are constructed such that they promote the presence of tunneling barriers Denchev et al. (2016); Mandra and Katzgraber (2018); Albash and Lidar (2018a). The competitive behavior that we observe in this work provides valuable insights into the potential advantages of quantum algorithms in the context of optimization problems.
We also show Dwave outperforms 1-round QAOA executed on hardware in terms of solution quality. Notably, QAOA demonstrates comparable runtime scaling with Dwave albeit with a significantly larger prefactor. 1-round QAOA on the simulator represents a solution quality lower but still close to Dwave up to the system sizes we have been able to simulate.
@@
Additionally, our experiments demonstrate that the solutions obtained through approximation on the ISG remain remarkably close to those derived from the original CFG formulation (up to system sizes that we have been able to check). This showcases the robustness and efficacy of our new introduced method in transferring CFG to ISG.
@@ II Problem formulation
$\min_{p_{i},i\in[C_{l}]}$ $\sum_{i\in[C_{l}]}c_{i}(p_{i})$ (1)
$-\Delta_{t}\sum_{t\in[T]}\left(\lambda_{t}^{b}\left[\sum_{i\in[C_{l}]}p_{it}\right]^{-}+\lambda_{t}^{s}\left[\sum_{i\in[C_{l}]}p_{it}\right]^{+}\right)$
[…]
We label the $N$ coalitions for which prosumers are metered individually $C^{0}_{i}$. Here, $\left[p_{it}\right]^{+}=\max\left\{0,p_{it}\right\}$ and $\left[p_{it}\right]^{-}=\min\left\{0,p_{it}\right\}.p_{it}$ is the prosumer output power during time interval $t$ and $p_{i}=\left(p_{i1},\ldots,p_{iT}\right)$ is the vector of output powers over the time horizon. $\Delta_{t}$ is the duration of each time interval. Each prosumer has a cost function $c_{i}\left(p_{i}\right)$ associated with the usage of their flexible devices and a constraint set $\mathcal{P}_{i}$ for their output power vector, which may be time-coupled. For the exact choice of $c_{i}\left(p_{i}\right)$ and constraint set $\mathcal{P}_{i}$, refer to Supplemental Material Sec. II.
@@
$*\displaystyle p_{i}^{\text{net* }}$ =$=\begin{cases}p_{i}^{*}\text{ from (1) optimized for $C_{l}$, }&\text{ if }i\in C_{l},\\ p_{i}^{*}\text{ from (1) optimized for $C^{0}_{i}$, }&\text{ otherwise }\end{cases}$
Here, indicates the solution to an optimization problem. $p_{i}^{\text{ind }*}$ ($p_{i}^{\text{net }*}$) represents the optimized power for the prosumer $i$ under individual (net) metering. $[N]={1,2,...N}$ and $\mathcal{O}^{opf*}$ is the objective function value for the following optimal power flow problem which the DSO can use to procure flexibility to manage losses and network constraints,
@@
The DNO decision problem is then to partition the prosumers into a set of net-metered coalitions with the goal of minimizing network management costs. This is formulated as a CFG problem, which can be solved using binary integer linear programming (BILP) Venkatesh et al. (2023b),
@@
s.t. $\sum_{j\in\left[2^{N}-1\right]}S_{il}x_{l}=1,\forall i\in[N],$
@@
Note that there are $12^{N}-$ 1 potential sub-coalitions of prosumers. $S_{il}=1$ if prosumer $i\in[N]$ belongs to coalition $C_{l}$ (i.e. $i\in C_{l}$ ), and $S_{il}=0$ otherwise. Binary decision variable $x_{l}=1$ indicates that $C_{l}$ is part of the selected partition, and the value function $v\left(C_{l}\right)$ gives the value of the coalition. This formulation is computationally intensive since the number of binary decision variables depends on the number of sub-coalitions, which increases exponentially with the number of prosumers (i.e. $12^{N}-$ 1).
This CFG formulation can be transformed into ISGs, where the value of a coalition can be expressed as the sum of pair-wise “joint utilities” between prosumers $w_{ij}\in\mathbb{R},i,j\in[N]$, i.e. $v^{\text{ISG}}\left(C_{l}\right)=\sum_{(i,j)\in C_{l}}w_{ij}$ with lower computational requirements. The value function nonlinearities mean this framework is not directly applicable to our CFG. However, we can make use of an ISG based on pairwise joint utilities $\tilde{w}_{ij}$ which approximate our CFG. The following quadratic programming problem can be used to find pairwise weights that minimize the mean squared error between the value functions and an approximate ISG,
@@
Starting with the grand coalition, an approximate solution to the ISG can be found by iteratively splitting coalitions in two, until no value-increasing bipartitions are available Venkatesh et al. (2023a). For a given iteration, let $C\in[N]$, be one of the optimal bipartitions that was found. The optimal bipartitions of $C$ at the next iteration can be found using the following quadratic unconstrained binary optimization (QUBO) with $N$ binary decision variables,
$\min_{{x}_{i},i\in C}$ $\sum_{i\in[N]}\sum_{j\in[N]\backslash i}\tilde{w}_{ij}x_{i}\left(1-x_{j}\right).$ (6)
@@
The schematic representation of the workflow is shown in Fig. 1.
@@ III Results
Figure 2: Benchmarking classical solvers against quantum solvers in energy coalition formation: (a) Solution quality of different solvers on ISG where the baseline is w.r.t. the solution determined with the exact solver on ISG. (b) The solution quality of different solvers on ISG w.r.t. the solution determined with IDP on the original CFG formulation. (c) Run time of different solvers on ISG as well as run time of improved dynamical programming (IDP) on CFG. Dwave and QB-solve curves are shown again in the inset to provide a closer examination of scaling. (d) QAOA circuit depth on quantum hardware for different numbers of prosumers (number of qubits). Results for each number of prosumers showcase averages across 20 realizations except for IBM hardware and random solver with $2^{12}$ where results are shown for a single problem instance per data point. The light blue histogram on the lower right shows the weight distribution after transferring CFG to ISG.
Before presenting the benchmarking results, it is essential first to elucidate the structure and complexity of our problem instances.
@@
Problem Structure: Our analysis reveals that the graphs resulting from the mapping of CFG to ISG are almost fully connected across all system sizes, with connectivity levels exceeding 95 percent. Moreover, the weight distribution includes both positive and negative values forming a Gaussian distribution centered closely around zero. The light blue histogram depicted in Fig. 2 lower right panel illustrates the weight distribution for the case of 18 prosumers averaging over 20 instances (different instances of the problem correspond to different initial power configurations for prosumers at the outset and varying export and import prices). This scenario, characterized by a fully connected graph featuring a mixture of positive and negative weights, represents a particularly challenging case McCormick et al. (2003). For further insights, referMcCormick et al. (2003), where they discuss the boundary between polynomially-solvable Max Cut and NP-Hard Max Cut instances when they are classified only on the basis of the sign pattern of the objective function coefficients.
Benchmark: Now, we elaborate on the outcome of an extensive performance evaluation, benchmarking an exact classical solver and QB-solve Booth et al. (2017) against quantum annealing on Dwave hardware (Dwave Advantage 4.1), and QAOA on both the simulator and IBMQ hardware (IBM-Osaka). QB-solve is a decomposing solver that uses the Tabu algorithm Palubeckis (2004); Glover (1986) for solving sub-problems. In addition to these solvers, we also include results obtained using a random sampler which is beneficial as it provides a baseline comparison to gauge the efficacy of the solvers. For a detailed overview of all the solvers utilized in our benchmarking process refer to Supplemental Material Sec. I.
@@
We benchmark the performances in terms of solution quality and running time (Fig. 2). The run time for all solvers encompasses any pre or post-processing time. The results for each number of prosumers showcase averages across 20 realizations except for IBM hardware and random solver with $2^{12}$ shots where results are shown for a single problem instance per data point.
Solution Quality: In Fig. 2 (a) and (b), we present the solution quality for both the transferred formulation, namely ISG, and the original formulation, CFG, respectively. For ISG, the solution quality, for up to 30 prosumers, is defined as the ratio of the value of the optimal coalition obtained from any solver on our ISG formulation to the value determined by the exact solver on our ISG formulation. In scenarios involving more than 30 prosumers, where the exact solution on ISG is unknown, the baseline is established by the optimal coalition obtained through QB-solve. Panel (b) shows the solution quality for the original CFG problem, where the solution quality is defined as the ratio of the optimal coalition value obtained from any solver on ISG to the baseline value determined by improved dynamic programming (IDP) on CFG. Further details about IDP can be found in Rahwan and Jennings (2008). Note that IDP is the only existing algorithm capable of finding an optimal solution for the CFG. Since IDP requires a considerable amount of time, typically takes several hours for 20 prosumers, thus limiting our ability to find solution quality for larger sizes.
Dwave and QB-solve exhibit comparable solution quality on both ISG and CFG. It is very encouraging to see that solutions obtained through approximation on the ISG remain remarkably close to those derived from the original CFG formulation. This confirms the efficacy of the mapping that we introduced in Eq. 5.
The 1-round QAOA executed on the simulator reveals a solution quality slightly lower, yet still close to that of Dwave and QB-solve. Enhancing the solution quality is still feasible by increasing the number of rounds. However, when executed on IBM hardware, the solution quality experiences a drastic decrease with system size. At times, it performs marginally better than a random sampler with the same number of shots, while at other times, it closely resembles the performance of a random solver. It is important to note that drawing a conclusive assessment would ideally involve averaging over a significant number of samples. However, such an analysis demands considerable resources, which we were unable to allocate.
Several factors contribute to the suboptimal performance of QAOA on hardware platforms, leaving room for improvement. Firstly, the original QAOA ansatz is inefficient and hardware-unfriendly for dense models. Recent studies have introduced new modified ansätze for QAOA that are more hardware-friendly Amaro et al. (2022); Yu et al. (2022). However, the effectiveness of these ansätze for fully connected models has yet to be thoroughly investigated. Secondly, additional efforts are necessary to transpile the circuit into a more hardware-friendly form Matsuo et al. (2023); Weidenfeller et al. (2022). As illustrated in Fig. 2 (d), the circuit depth increases quadratically with the number of qubits. The existing transpiling methods may improve the efficiency of QAOA but they still highlight further obstacles to improve to make QAOA competitiveMatsuo et al. (2023); Weidenfeller et al. (2022). Moreover, here we just applied error mitigation in the readout, implementing additional mitigation schemes is required for QAOA Sack and Egger (2024).
@@
Run Time: Fig. 2 (c), shows the run time of different solvers. For a closer examination of scaling, we have included Dwave and QB-solve run time separately again in the inset. Dwave and QB-solve demonstrate competitive runtime however up to 45 prosumers it is hard to see a clear scaling. The exact classical solver on ISG and CFG, due to their computational complexity, experience a noticeable increase in runtime that become impractical beyond 30 prosumers. The runtime of QAOA on the simulator exhibits polynomial scaling, though sub-optimal. Interestingly, the scaling behavior of QAOA on Hardware is close to Dwave and QB-solve, albeit with a significantly larger prefactor. This can be attributed to iteratively updating variational parameters and running the algorithm on hardware for each parameter update. Initializing variational parameters based on existing heuristic schemes Zhou et al. (2020); Akshay et al. (2021) or learning them via machine learning techniques Montanez-Barrera et al. (2024); Mohseni et al. (2023) could significantly reduce this prefactor.
@@
Figure 3: Benchmarking classical solvers against Dwave in energy coalition formation for randomly generated problems (sourced from the same distribution as the real use case problems namely a Gaussian distribution with a width of 2). (a) The solution quality (Average value per prosumer) versus the number of prosumers. (b) Run time versus number of prosumers. Results for each number of prosumers showcase averages across 20 realizations. (c) The number of physical qubits versus logical qubits (number of prosumers) on Dwave hardware for the first split of the grand coalition into two partitions.

### v2

# Demonstrating Quantum Scaling Advantage in Approximate Optimization for Energy Coalition Formation with 100+ Agents

@@ Abstract
The formation of energy communities is pivotal for advancing decentralized and sustainable energy management. Within this context, Coalition Structure Generation (CSG) emerges as a promising framework. The complexity of CSG grows rapidly with the number of agents, making classical solvers impractical for even moderate sizes. This suggests CSG as an ideal candidate for benchmarking quantum algorithms against classical ones. Facing ongoing challenges in attaining computational quantum advantage for exact optimization, we pivot our focus to benchmarking quantum and classical solvers for approximate optimization. Approximate optimization is particularly critical for industrial use cases requiring real-time optimization, where finding high-quality solutions quickly is often more valuable than achieving exact solutions more slowly. Our findings indicate that quantum annealing (QA) on DWave can achieve solutions of comparable quality to our best classical solver, but with more favorable runtime scaling, showcasing an advantage. This advantage is observed when compared to solvers, such as Tabu search, simulated annealing, and the state-of-the-art solver Gurobi in finding approximate solutions for energy community formation involving over 100 agents. DWave also surpasses 1-round QAOA on IBM hardware. Our findings represent the largest benchmark of quantum approximate optimizations for a real-world dense model beyond the hardware’s native topology, where D-Wave demonstrates a scaling advantage.
@@ I Introduction
To explore the potential of quantum computing in optimization problems it is encouraging to benchmark quantum solvers against classical solvers in tackling problems that are sufficiently hard and exceed the capabilities of state-of-the-art classical solvers. This exploration becomes even more intriguing when problems get already hard in intermediate sizes where existing quantum hardware can be effectively leveraged. It is also crucial that the problem can be formulated compatible with existing quantum algorithms and hardware Abbas et al. (2023). As we explain next, energy coalition formation applying the Coalition Structure Generation (CSG) game stands out as an exemplary candidate for such benchmarking, given its computational complexity and profound real-world significance within the industry. The formation of energy communities represents a significant step towards decentralized and sustainable energy management, offering opportunities for local empowerment, renewable energy adoption, and enhanced resilience in energy systems Tushar et al. (2020); Luo and Liu (2021); Blenninger et al. (2024, 2024); Nüßlein et al. (2024); Bucher et al. (2024).
@@
CSG provides a framework for understanding how agents, can form coalitions denoted by $C_{l}$ (with $l$ being the index of the coalition) to optimize their collective benefits Bandeiras et al. (2023); Moafi et al. (2023); Han et al. (2018); Rahwan et al. (2015). A coalition within this context can be characterized by a characteristic function, typically denoted as $v(C_{l})$, which assigns values to different coalitions. The value function provides a measure of the cooperative advantage or payoff that the agents in the coalition can attain by working together, as opposed to operating independently Bachrach et al. (2013); Deng and Papadimitriou (1994).
@@
The task of creating coalition structures is highly challenging, especially given the presence of $n$ agents, leading to a staggering $O(n^{n})$ possible partitions. Currently, the only algorithm capable of finding an optimal solution is the Dynamic Programming algorithm, achieving this in $O(3^{n})$ time Rahwan and Jennings (2008). This approach is limited to systems consisting of a few tens of agents ( i.e. $n<$ 25). CSG can be transformed into an approximately equivalent game represented using the induced subgraph game (ISG) Bistaffa et al. (2022). In this representation, nodes correspond to agents, and coalition values are reflected in pairwise interactions within a weighted graph, such that $v\left(C_{l}\right)=\sum_{(i,j)\in C_{l}}w_{ij}$ Deng and Papadimitriou (1994); Bachrach et al. (2013). The transformation of CSG to ISG allows the problem to be recast as quadratic unconstrained binary optimization (QUBO), enabling more computationally efficient approaches. Nonetheless, it is crucial to note that the problem remains NP-complete Bachrach et al. (2013) such that leading classical solvers, such as DyCE Bistaffa et al. (2014) and CFSS Bistaffa et al. (2014) are limited in scalability and memory requirements. For instance, DyCE Bistaffa et al. (2014), which employs dynamic programming, is restricted to systems with around 25 agents due to its exponential memory complexity ($O\left(2^{n}\right)$) and lacks an anytime approach. CFSS can handle larger but shallow problems, yet still explores all possible solutions for complete graphs with a complexity of $O(n^{n})$.
@@
Figure 1: (a) Schematic representation of the workflow: The Distribution System Operator (DSO) strategically evaluates the potential of prosumer communities, aiming to minimize network management costs through optimized power flow analysis. The DSO’s decision-making involves partitioning prosumers into net-metered coalitions, treating it as a Coalition Structure Generation (CSG) problem to optimize overall network efficiency. We transform the CSG formulation into the Induced Subgraph Game (ISG). An approximate solution to the ISG is found through iterative splitting coalitions to bipartitions until no value-increasing bipartitions remain. In the benchmarking phase, the efficiency of both quantum and classical solvers is assessed in the iterative splitting process.
In this work, we propose a novel formulation of energy coalition formation that can be encapsulated within the framework of CSG. We first introduce a novel method for transforming the CSG formulation into an ISG. Building on this transformation, we apply the approximation introduced in Venkatesh et al. (2023), under which our problem reduces to an iterative weighted minimum cut on the ISG (formulated as a QUBO). The complexity of finding an optimal solution for this reduced formulation remains $O(n^{3})$. We show a solution where its optimality is not proven can be found in polynomial time.
Our performance evaluation involves benchmarking quantum annealing on DWave hardware and the Quantum Approximation Optimization Algorithm (QAOA) on IBMQ hardware against classical solvers such as Gurobi Gurobi Optimization, LLC (2024), Tabu search Glover (1986), simulated annealing Kirkpatrick et al. (1983), and QB-solve Booth et al. (2017)) to find approximate solutions. We provide preliminary empirical evidence suggesting a potential scaling advantage of DWave for a dense, real-world problem in approximate optimization. A recent study Bauza and Lidar (2025) applying quantum annealing correction demonstrates a scaling advantage for approximate optimization against parallel tempering with isoenergetic cluster moves. However, that work is limited by the problem’s compatibility with the hardware’s connectivity and is focused on a 2D model, which restricts its range of applications. We also show Dwave outperforms 1-round QAOA on IBM Hardware.
@@
Additionally, our experiments demonstrate that the solutions obtained through approximation on the ISG remain remarkably close to those derived from the original CSG formulation (up to system sizes that we have been able to check). This showcases the robustness and efficacy of our new introduced method in transferring CSG to ISG.
@@ II Problem formulation
$\min_{p_{i},i\in[C_{l}]}$ $\left(\sum_{i\in[C_{l}]}c_{i}(p_{i})\right$. (1)
$\left.-\Delta_{t}\sum_{t\in[T]}\left(\lambda_{t}^{b}\left[\sum_{i\in[C_{l}]}p_{it}\right]^{-}+\lambda_{t}^{s}\left[\sum_{i\in[C_{l}]}p_{it}\right]^{+}\right)\right)$
$\text{s.t. }p_{i}\in\mathcal{P}_{i},\forall i\in[C_{l}].$
$c_{i}\left(p_{i}\right)$ presents each prosumer’s cost function associated with the usage of their flexible devices and a constraint set $\mathcal{P}_{i}$ for their output power vector, which may be time-coupled. For the exact choice of $c_{i}\left(p_{i}\right)$ and constraint set $\mathcal{P}_{i}$, refer to Supplemental Material Sec. II. We label the $N$ coalitions for which prosumers are metered individually $C^{0}_{i}$. We remind that $C_{i}$ refers to $i$-th colaition. $p_{i}=\left(p_{i1},\ldots,p_{iT}\right)$ is the vector of output powers over the time horizon and $p_{it}$ is the prosumer output power during time interval $t$. $\left[p_{it}\right]^{+}=\max\left\{0,p_{it}\right\}$ and $\left[p_{it}\right]^{-}=\min\left\{0,p_{it}\right\}$. $\Delta_{t}$ is the duration of each time interval.
@@
[…]
Here, indicates the solution to an optimization problem. $p_{i}^{\text{ind }*}$ ($p_{i}^{\text{net }*}$) represents the optimized power for the prosumer $i$ under individual (net) metering. $N]=\{1,2,...N\}$ and $\mathcal{O}^{opf*}$ is the objective function value for the following optimal power flow problem which the DSO can use to procure flexibility to manage losses and network constraints,
@@
The DSO decision problem is then to partition the prosumers into a set of net-metered coalitions with the goal of minimizing network management costs. This is formulated as a CSG problem, which can be solved using binary integer linear programming (BILP) Venkatesh et al. (2022),
@@
s.t. $\sum_{i\in\left[2^{N}-1\right]}S_{il}x_{l}=1,\forall i\in[N],$
@@
Note that there are $2^{N}-1$ potential sub-coalitions of prosumers. $S_{il}=1$ if prosumer $i\in[N]$ belongs to coalition $C_{l}$ (i.e. $i\in C_{l}$ ), and $S_{il}=0$ otherwise. Binary decision variable $x_{l}=1$ indicates that $C_{l}$ is part of the selected partition, and the value function $v\left(C_{l}\right)$ gives the value of the coalition. This formulation is computationally intensive since the number of binary decision variables depends on the number of sub-coalitions, which increases exponentially with the number of prosumers (i.e. $2^{N}-1$).
This CSG formulation can be transformed into ISGs, where the value of a coalition can be expressed as the sum of pair-wise “joint utilities” between prosumers $w_{ij}\in\mathbb{R},i,j\in[N]$, i.e. $v^{\text{ISG}}\left(C_{l}\right)=\sum_{(i,j)\in C_{l}}w_{ij}$ with lower computational requirements. The value function nonlinearities mean this framework is not directly applicable to our CSG. However, we can make use of an ISG based on pairwise joint utilities $\tilde{w}_{ij}$ which approximate our CSG. The following quadratic programming problem can be used to find pairwise weights that minimize the mean squared error between the value functions and an approximate ISG,
@@
Starting with the grand coalition, an approximate solution to the ISG can be found by iteratively splitting coalitions in two, until no value-increasing bipartitions are available Venkatesh et al. (2023). For a given iteration, let $C$, be one of the optimal bipartitions that was found for our $N$ prosumers. The optimal bipartitions of $C$ at the next iteration can be found using the following quadratic unconstrained binary optimization (QUBO) with $N$ binary decision variables $x_{i}$,
[…]
@@
Here the binary vector $\mathbf{x}=\left[x_{1},x_{2},\ldots,x_{N}\right]$ for $N$ prosumers encodes the bipartition: prosumers with the same value of $x_{i}$ belong to the same subcoalition. For example, $\mathbf{x}=[1,0,0,1]$ indicates that prosumers 1 and 4 form one subcoalition, while prosumers 2 and 3 form the other.
The schematic representation of the full workflow is shown in Fig. 1.
@@ III Results
Figure 2: Benchmarking classical solvers against quantum solvers in energy coalition formation: (a) Solution quality of different solvers on ISG formulation where the baseline is w.r.t. the solution determined with the Gurobi on the same formulation. (b) The solution quality of different solvers on ISG formulation w.r.t. the solution determined with IDP on the original CSG formulation. (c) Run time of different solvers on ISG formulations as well as run time of improved dynamical programming (IDP) on CSG formulation. (d) QAOA circuit depth on quantum hardware for different numbers of prosumers (number of qubits). Note that results for each number of prosumers showcase averages across 20 realizations except for IBM hardware and random solver with $2^{12}$ shots where results are shown for a single problem instance per data point. The light blue histogram on the lower right shows the weight distribution after transferring CSG to ISG.
In this section, we elaborate on the outcome of our performance evaluation, benchmarking an exact classical solver, the commercial state-of-the-art classical solver Gurobi, and QB-solve Booth et al. (2017) against quantum annealing on DWave hardware (DWave Advantage 4.1), and 1-round QAOA on IBMQ hardware (IBM-Osaka). For a detailed overview of all the solvers utilized in our benchmarking, refer to Supplemental Material Sec. I. Before presenting the benchmarking results, it is essential to elucidate the structure and complexity of the problem instances generated from our use case as follows.
@@
Problem complexity: Our analysis reveals that the graphs resulting from the mapping of CSG to ISG are very dense across all system sizes, with connectivity levels exceeding 90 percent. Moreover, the weight distribution includes both positive and negative values forming a Gaussian distribution centered closely around zero. The light blue histogram depicted in Fig. 2 lower right panel illustrates the weight distribution for the case of 18 prosumers averaging over 20 instances (different instances of the problem correspond to different initial power configurations for prosumers at the outset and varying export and import prices). This scenario, characterized by a dense graph featuring a mixture of positive and negative weights, represents a particularly challenging case McCormick et al. (2003). For further insights, refer toMcCormick et al. (2003), where they discuss the boundary between polynomially-solvable Max Cut and NP-Hard Max Cut instances when they are classified only on the basis of the sign pattern of the objective function coefficients.
Solver compatibility: When benchmarking algorithms, selecting the appropriate formulation for each solver is crucial to ensure their optimal performance. Our problem formulation here namely the QUBO formulation is well-suited to all the solvers we are applying in this work.
@@
Benchmark metrics: We benchmark the performances in terms of solution quality and running time (Fig. 2). The runtime for all solvers includes also any classical processing time, such as the time spent optimizing variational parameters. The results for each number of prosumers showcase averages across 20 realizations except for IBM hardware and random solver with $2^{12}$ shots where results are shown for a single problem instance per data point.
Solution quality: In Fig. 2(a), we present the solution quality defined as the ratio of the optimal coalition values obtained by any solver on ISG formulation to the values found by Gurobi on the same formulation. Up to this size, Gurobi is still able to prove the optimality of the solutions found. For the prosumer sizes studied here (up to 45), both DWave and our classical solver consistently match Gurobi’s solution. In contrast, the 1-round QAOA exhibits a significant decline in solution quality as system size increases and its performance is indistinguishable from a random solver. In some cases, QAOA performs slightly better than a random sampler but mostly its performance is indistinguishable from a random solver.
Here QAOA is enhanced with Fire Opal’s error suppression pipeline, excluding any readout mitigation and postprocessing. This incorporates automated error-suppression techniques, such as layout optimization Hartnett et al. (2024), dynamical decoupling for crosstalk and dephasing suppression Ezzell et al. (2023), and AI-driven gate waveform optimization.
Several factors contribute to the suboptimal performance of QAOA on IBM hardware. The most significant one is that the traditional QAOA ansatz is inefficient and not hardware-friendly for dense problems like ours. In Fig. 2 (d), we compare the circuit depth applying the Fire Opal transpiler and the default Qiskit transpiler, showing that circuit depth increases quadratically with the number of qubits using Qiskit’s default. While Fire Opal’s transpilation significantly reduces the circuit depth and improves scaling, it still remains insufficient to make QAOA competitive for dense problems, given current gate fidelities Matsuo et al. (2023); Weidenfeller et al. (2022). Recent studies have proposed hardware-friendly modifications to the QAOA ansatz Amaro et al. (2022); Yu et al. (2022), but their effectiveness for dense models remains elusive.
We also investigate the closeness of the solution derived from the ISG transformation to that derived from the original CSG formulation in Panel (b). In this case, the solution quality is defined as the ratio of the optimal coalition value obtained from any solver on ISG to the exact solution of the original CSG formulation determined by improved dynamic programming (IDP) Rahwan and Jennings (2008). IDP is the only existing algorithm capable of finding an optimal solution for the CSG formulation. However, the runtime scales as $O(n^{3})$, allowing us to calculate the solution quality for up to 20 prosumers. It is very encouraging to see that solutions obtained through approximation on the ISG remain remarkably close to those derived from the original CSG formulation. This confirms the efficacy of the mapping that we introduced in Eq. 5.
Run Time: Fig. 2 (c), shows the run time of different solvers. DWave and QB-Solve demonstrate competitive runtimes, but clear scaling is difficult to observe with up to 45 prosumers. For Gurobi, we report two runtimes: the cyan line represents the time taken to find a solution that no longer improves over time, though its optimality is not proven. The brown line shows the total runtime, including the time spent proving the solution’s optimality. As the number of prosumers increases, it becomes increasingly challenging for Gurobi to confirm optimality. For instances with more than 50 prosumers, Gurobi is unable to prove optimality even after several hours of computation. IDP on CSG also experiences a noticeable increase in runtime as the problem size grows.
@@
Interestingly, the scaling behavior of QAOA on Hardware is close to DWave and QB-solve, albeit with a significantly larger prefactor. This can be attributed to iteratively updating variational parameters and running the algorithm on hardware for each parameter update. Initializing variational parameters based on existing heuristic schemes Zhou et al. (2020); Akshay et al. (2021) or learning them via machine learning techniques Montanez-Barrera et al. (2024); Mohseni et al. (2023a, b) could potentially reduce this prefactor to some extent.
To compare our solvers effectively, it is essential to consider both runtime and solution quality together. For problem sizes up to 45 prosumers, both D-Wave and our classical solver consistently achieve matching solutions; however, runtime comparisons at this scale do not provide reliable insights into scaling behavior. Our problem instances become difficult to solve beyond 50 prosumers, where even Gurobi struggles to find optimal solutions. Therefore, drawing firm conclusions about performance requires exploring larger instances, which we address next.
@@
Figure 3: Benchmarking classical solvers against DWave in energy coalition formation for randomly generated problem instances. (a) The solution quality versus the number of prosumers. (b) Run time versus number of prosumers. (c) The number of physical qubits versus logical qubits (number of prosumers) on DWave hardware for the first split of the grand coalition into two partitions. Note that results for each number of prosumers showcase averages across 20 realizations. (d) Energy distribution sampled from DWave, QAOA, and a random sampler for the first split of the grand coalition to two partitions for a randomly generated instance with 100 prosumers (100 qubits). The dashed vertical line shows the solution found by Gurobi.